OpenStudy (anonymous):

find y such that the distance between the points is 8. (5,1) and (5,y)

OpenStudy (amistre64):

1+8 = ?

OpenStudy (amistre64):

or 1-8 as well

myininaya (myininaya):

\[8=\sqrt{(5-5)^2+(1-y)^2}\] \[8=1-y\] \[8-1=-y\] y=-7

myininaya (myininaya):

or -8=1-y -8-1=-y y=9

OpenStudy (amistre64):

distance formula for measuring on vertical line; cute :)

OpenStudy (anonymous):

oh lol thx

OpenStudy (anonymous):

x coordinates are same so this is a "vertical" line through x = 5.

myininaya (myininaya):

lol